OpenStudy (anonymous):

what is the speed of a 0.145 kg baseball if its kinetic energy is 109 J?

OpenStudy (s3a):

E_k = 1/2 * m * v^2 109 J = 1/2 * (0.145kg) * v^2 (109 J)(2)/(0.145 kg) = v^2 v = sqrt((109 J)(2)/(0.145 kg)) v = 38.77 m/s